OpenStudy (anonymous):

jen is 13 years younger than andre. the sum of their ages in years in 137. what is andre's age in years?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

just subtract the two

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

I don't think that's right... when you subtract them you get 124..

OpenStudy (anonymous):

Andre = x Jen = y y = x -13 x + y = 137 Solve the system

OpenStudy (anonymous):

And then you have double Jen's age. This is basically 2x+13=134=2y-13, and you can solve either or both sides.

OpenStudy (anonymous):

Knock I dont know how to solve the system of equations. could you shoe me how?

OpenStudy (anonymous):

show**

OpenStudy (anonymous):

Sure! There are different ways, I think the faster here is: Andre = x Jen = y 1)y = x -13 2)x + y = 137 From 2) y = 137 - x, now substitute y in 1) and 137 - x = x - 13; 150 = 2x; x = 150 / 2; x = 75 (Andre) With that value you go to 1) and y = 75 - 13 = 62; y = 62 (Jen)

OpenStudy (anonymous):

Thank you soooooo much!!!!! =]

OpenStudy (anonymous):

Yw (:

OpenStudy (anonymous):

You kinda look like Robert Downey Jr.